OpenStudy (anonymous):

Which of these ordered pairs is a solution to the linear inequality y – 4x ≤ –6? A. (–2, 4) B. (1, –2) C. (1, 3) D. (2, 3)

OpenStudy (anonymous):

Plug in each of your points and see if what results is true. For example, take the first one and put -2 and 4 in for x and y. x=-2, y=4 \[4-4(-2)\le -6\] \[4+8\le -6\] \[12 \le -14\] That's not true, 12 is greater than -14, so move on to the next ordered pair. Hope that helps :)

OpenStudy (anonymous):

The ordered pair that results in a true inequality is your answer.

OpenStudy (anonymous):

Oops idk why I put -14, it should be -6

OpenStudy (anonymous):

so is A my answer

OpenStudy (anonymous):

No, you have to keep trying your options til one results in a true inequality. I don't want to do all the work for you :)

OpenStudy (anonymous):

I'm trying but all of this is kinda confusing to me

OpenStudy (anonymous):

I'm thinking its C

OpenStudy (anonymous):

Well, you do basically the same process with the other answers: Just plug in the x and y values for x and y in your inequality. For B, x=1 and y=-2 So replace x with 1 and y with -2: \[-2-4(1)\le -6\] And then you just solve. -4*1=-4. -2-4=-6. So it just so happens that B is your answer because -6 is equal to -6 (the sign says the solution is greater than OR equal to)

OpenStudy (anonymous):

I hope that makes sense :)

OpenStudy (anonymous):

C wouldn't be your answer because you end up with: \[-1\le -6\] but that's not true because -1 is greater than -6, not less than -6.

OpenStudy (anonymous):

oh okay I understand what I was doing wrong thanks

OpenStudy (anonymous):

Yay! :) No problem.

OpenStudy (anonymous):

how do I graph x < 2.? and 3y – 5x > –6

OpenStudy (anonymous):

For x<2, you want to make your dot 2 from the origin on the x-axis, like this:  then make a vertical, DOTTED line, because it is <

OpenStudy (anonymous):

would there be shading to the left or right.? and thanks

OpenStudy (anonymous):

For the other one, we first need to simplify it like this: 3y-5x>-6 3y-5x+5x>-6+5x 3y>5x-6 3y/3>5x/3-6/3 y>5/3x-2 So you'll start by making your first dot down 2 on the y-axis, like this: then you go up five, over 3 to make the second dot, and this again with be a dotted line.

OpenStudy (anonymous):

Oh sorry :3 x is LESS THAN so you shade on the left.

OpenStudy (anonymous):

for the next one is that also shaded to left or is it right

OpenStudy (anonymous):

It is GREATER THAN so that means it's shaded on the right :)

OpenStudy (anonymous):

sorry I suck at Algebra... and okay thats what I thought and I just wanted to make sure thanks again for all the help

OpenStudy (anonymous):

It's okay, I understand. I hate algebra, it sucks xP But after you've done it for a long enough time, it starts to make some sense. No problem! Good luck :)